(51 – 1) is 37. Use the Principle of Mathematical Induction to show 1 +5 +52 + ... 5n-1 = true for all natural numbers n.
Answer #1

Let the given Statement is pan) P(n) = 1+ 5 + 5² - tshi 1 (5-1) 4 For n=1 P(1) = 1 = 1 (5-1) 4 1 = 1x4 4 1 1 which is true As